OpenStudy (anonymous):

help...This is my little bothers homework and he needs help....Its estimate products of fractions and here is the question...Tara would like to finish 2/5 of her cook by next Friday if the book has 203 pages about how many pages does she need to read? please give step by step answers...thanks

OpenStudy (pixiedust1):

Okay! So there are 203 pages. Tara wants to read 2/5 of it. Let's first figure out how many pages in every 1/5. Since this book is being divided into 5ths, lets divide 203/5=40.6 So, every 1/5 of the book is 40.6 pages. But the question asked for every 2/5. So let's add another 1/5. 1/5+1/5=2/5 40.6+40.6=81.2 pages!

OpenStudy (pixiedust1):

If you need to round, you would round down to 81

OpenStudy (anonymous):

This is my little brother talking and me "thanks"

OpenStudy (pixiedust1):

Your welcome, Izzy's little brother! Got any other questions?

OpenStudy (anonymous):

he said yes lol

OpenStudy (pixiedust1):

lol! :D

OpenStudy (pixiedust1):

Well I'd be happy to help :)

OpenStudy (anonymous):

Okay so here is the question and you should ask him questions as you go step by step to he him understand... Seattle Washington,received rain on 7/10 of the days in a recent month if the pattern continues, about how many days would it not rain in 90 days?

OpenStudy (pixiedust1):

Okay! Hello Izzy's little brother! If it rains for 7/10 days, that means it DOESN'T"T rain for 3/10 days So, this question basically wants us to know about 3/10. Understand so far?

OpenStudy (anonymous):

he said oh okay and his name is trey lol

OpenStudy (pixiedust1):

Okay Trey! (That kinda rhymes!) Trey, what do you think the next step would be?

OpenStudy (pixiedust1):

If it doesn't rain for 3/10 days, then we are basically finding 3/10 of 90. How would we do that?

OpenStudy (anonymous):

we could turn the fraction into a decimal?

OpenStudy (anonymous):

then divide 90 by that number?

OpenStudy (pixiedust1):

Close! We would actually divide by the "10" in the 3/10. So, 90/10=9 So, we now know that 1/10= 9 days So, if we know how many days is 1/10, how do you think we would find out 3/10

OpenStudy (anonymous):

9 times 9 times 9

OpenStudy (anonymous):

27

OpenStudy (pixiedust1):

Once again, pretty close! 9*9*9=729, which is way more days than 90! But we would do, 9*3 (remember, 3/10), and yes, 9*3=27! :D

OpenStudy (pixiedust1):

So, we now know that it would not rain 27 days out of 90 in Seattle ;)

OpenStudy (anonymous):

oh I was thinking 9 times 3 in my head so i guess i kind of got it

OpenStudy (pixiedust1):

Oh good :D

OpenStudy (anonymous):

so 27 is the answer

OpenStudy (pixiedust1):

You bet! :)

OpenStudy (anonymous):

wow this was easier then i thought

OpenStudy (pixiedust1):

Don't forget to put a label... 27 days without rain in Seattle, Washington Some teachers really push for the label, but others aren't so strict with it.

OpenStudy (anonymous):

okay

OpenStudy (pixiedust1):

If you have any other questions, just let me know! If not, have a blessed day, Izzy and Trey!

OpenStudy (anonymous):

yeah were good for now you too thanks